# Thermoelectric properties of defective scandium nitride nanostructures

### v1

@@ Abstract
Transition-metal nitrides (TMNs) are currently being studied for potential applications in energy conversion. In this work, we used the Landauer approach to relate the various effects contributing to the thermoelectric efficiency of scandium nitride (ScN) to their microscopic origins. We model the impact of electronic and structural modifications induced by oxygen impurities and spatial vacancies on electronic transport in ScN nanostructures. Taking advantage of the results of our calculations, we propose a theoretical interpretation of recent experimental results revealing a strong dependence of the thermoelectric properties of ScN thin films on procedural variations during fabrication. The thermoelectric properties of ScN are decisively influenced by structural and electronic factors arising from defects or impurities. Our findings highlight the potential of this theoretical approach in studying thermoelectricity and uncovering future strategies to improve thermoelectric efficiency.
@@ I Introduction
Appropriate preparation techniques for ScN thin films can also significantly reduce their thermal conductivities, as observed with lithium (Li) ions irradiation during plasma-assisted molecular beam epitaxy [22], or by implanting argon (Ar) atoms into ScN thin films during reactive magnetron sputtering deposition using an ion beam implantation technique [23]. These procedures lead to a strong decrease in the in-plane thermal conductivity of the resulting films, reaching respectively 7 and 3 $\mathrm{W/(m\cdot K)}$ at RT. Recently, a similar reduction has been reported on ScN thin films prepared by reactive magnetron sputtering, with periodic interruptions of the cathodic potential during the deposition procedure [24]. However, in addition to their impact on thermal transport, these techniques also affect electronic transport, leading to opposing effects—regarding thermoelectric efficiency—between the Seebeck coefficient and electrical conductivity. Specifically, they induce a significant increase—in absolute terms—of the Seebeck coefficient while causing a strong reduction in electrical conductivity. The total effect on the $zT$ value is slightly positive, with an increase in the thermoelectric figure of merit primarily driven by the substantial reduction in thermal conductivity, despite a decrease in the thermoelectric power factor [22, 23, 24].
Theoretical studies have drawn important relations between the transport properties of ScN and its structural defects. Impurities and lattice vacancies have been identified as determinant aspects for thermoelectricity in ScN [25, 26, 27, 28, 29]. Understanding these microscopic regulation mechanisms is the key to access further optimization techniques able to selectively tune the individual parameters without negatively affecting others.
In this work, we want to explore the underlying structure of the observed experimental responses [22, 23, 24]. We aim to do so by modeling the electronic transport in ScN nanowires (NW) and associating it with structural lattice defects, whose formation in the ScN lattice can be hypothesized under the described experimental conditions [22, 23, 24]. We find general conclusions applicable to ScN lattices regardless of their dimensions. Specifically, we study point defects such as the inclusion of oxygen (O) impurities and atomic vacancies in place of the lattice nitrogen (N) atoms. Furthermore, we examine the effects of association of multiple defects in contiguous lattice positions and the effects of structural planar gliding defects.
@@ II.1 The Landauer approach to quantum transport
$\Sigma_{\{s,d\}}(E)$ and $\Sigma_{\{s,d\}}^{\dagger}(E)$ are the retarded and advanced self-energies of the leads (source or drain), which contain the retarded and advanced Green’s functions of the leads and the coupling matrices between the central conductor and the leads. Their rigorous expression can be found e.g., in Refs. 36, 33, 37. The retarded and advanced Green’s functions of the central conductor, $\mathcal{G}_{c}^{r}(E)$ and $\mathcal{G}_{c}^{a}(E)$, describe the propagation of the electrons forwards and backwards in time in this section of the system. These can be straightforwardly calculated once knowing the real-space Hamiltonian of the system and the Green’s functions of the semi-infinite leads [36, 37, 33].
In this work, we calculate the Bloch space Hamiltonians of the systems using plane waves Density Functional Theory [38] and the real-space Hamiltonians with a subsequent projection and filtering procedure [39]. The Green’s functions of the semi-infinite leads are then calculated dividing the real-space representation of the system into a series of principal layers orthogonal to the transport direction and then using the iterative Lopez-Sancho et al. procedure [40, 35].
@@ II.2 Structural models
As mentioned, ScN has a face-centered cubic (fcc) rocksalt structure with a calculated lattice constant of 4.530 Å. The primitive cell contains 2 atoms per cell, while it is possible to define a conventional simple cubic (sc) cell using a larger crystal unit, containing 8 atoms per cell. The axes of the conventional sc cell are aligned along the Sc–N bonds directions. Using periodic boundary conditions (PBC), we built the NW models along one of the Sc–N direction. The NWs were constructed by inserting sufficient space in the directions orthogonal to their length and replicating half of the surface layers of the obtained NW on the opposite sides (these additional layers are represented by +s in the symbolism that we introduce to denote the NWs’geometries). This operation was done to make the cross-section of the NWs centrosymmetric. Two different periodicities were employed, corresponding to either one or three conventional cells along the NW direction, and different cross-section areas: (1x)1x1+s, (1x)3x3+s, (1x)5x5+s, (3x)3x3+s. These symbols describe the models’ structures: in parentheses the length of the periodic unit—expressed in conventional cells—then the cross-sectional area, and the symbol +s for additional layers.
@@
Figure 2: Schematic representations of (a) a glide-type defect, (b) a point defect, and (c) a combined configuration of a glide-type and a point defect.
The possible structural defects that may arise in the ScN lattice can be classified according to their symmetry into point defects and glide-type defects. Combinations of these two types of defects can also occur in spatially adjacent positions (Figure 2).
The structure of ScN allows only one type of symmetry for glide-type defects, with a displacement parallel to the Sc–N bond direction. We modeled glide-type defects associated with point defects in the (1x)5x5+s structures, after obtaining the relaxed structures with point defects, by replacing an adjacent relaxed atomic layer, perpendicular to the NW length, with a layer from the unrelaxed structure, centered along the NW axis.
In contrast, several types of point-type defects can be accounted: vacancies and substitutions of either Sc or N atoms, or clusters of multiple such defects. These defects can be ranked based on their relevance by calculating their formation enthalpy, revealing that substitutional N defects are the most energetically favorable and thus the most common in experimental samples [41].
Concerning the chemical nature of defects, under magnetron sputtering deposition or molecular-beam epitaxy experimental conditions [19, 20, 22, 23, 24], the presence of oxygen impurities is likely unavoidable, even in ultra-high-vacuum (UHV) systems [42, 43]. We studied substitutional defects where N sites are replaced either by structural vacancies or by oxygen atoms, and association of these defects with contiguous point or glide-type defects. We coherently compared each case to models of pure ScN nanowires of the same size.
@@ II.3 Detailed description
The atomic positions in systems with structural point defects were obtained through structural relaxation calculations performed using the Broyden–Fletcher–Goldfarb–Shanno (BFGS) algorithm [48]. These structural relaxations were carried out on bulk systems, constructed with PBC. We introduced the structural defect at the center of supercells of size 3x3x3 and 1x5x5 conventional cells. For the structural relaxation calculations, the eigenvalues were calculated on uniform 3x3x3 k-points grids spanning the whole first Brillouin zone of the systems; we employed a PW basis set defined by a cutoff energy of $40\,\text{Ry}$; the electronic charge density was recalculated after each diagonalization using a larger set of plane waves, defined by a cutoff energy of $320\,\text{Ry}$ before proceeding to the next step in the self-consistent field (SCF) cycle. The convergence thresholds were set to $10^{-6}$ a.u. for the total energy and $10^{-5}$ a.u. for the forces.
@@ III Results and discussion
In Figure 3, we present the calculated electronic band structures of pure ScN represented along high-symmetry lines of the primitive (p) face-centered cubic (fcc) lattice—containing two atoms per unit cell—and the conventional (c) simple cubic (sc) lattice—containing eight atoms per cell. The indirect band gap $\Gamma\text{--}X$]p observed in the fcc representation appears as a direct gap at the $\Gamma$ point in the sc representation, due to band folding at the $\Gamma$ point. Theoretical literature reports an indirect $\Gamma\text{--}X$]p band gap of $\sim 0.9$ eV for pure ScN, calculated using quasiparticle corrections [50, 51], exact-exchange approach [7] or hybrid functionals [52, 8]. These theoretical values are in good agreement with experimental literature [12, 53, 7]. This indirect band gap, however, is not captured by DFT, due to the well-known underestimation of the fundamental gap. In either the local density approximation (LDA) or GGA (Figure 3), it is predicted to be zero [54]. Apart from this known issue, the electronic structure computed using DFT does not exhibit significant topological or analytic variations in the shape of the bands near to the Fermi energy when compared to those obtained using quasiparticle corrections, exact exchange or hybrid functionals [50, 25].
@@
In Figure 4, we compare the electronic band structures of the defected (1×)5×5+s systems with central oxygen atoms and vacancies. While the presence of oxygen only leads to an $n$-type doping effect, with a rigid shift of the Fermi level without significantly distorting the band topology, the system with vacancies, in addition to a similar effect, exhibits also a strong distortion in the shape of a single band—whose real-space representation is shown in Figure 5—to the point that it crosses the Fermi level. This, in turn, affects the profile of the electronic transmission function (Figure 6), with subsequent effect on the Seebeck coefficient.
@@
Figure 5: A section of the 155v NW system. In gray: real-space representation at the $\Gamma$-point of the wavefunction associated with the electronic state whose band connects the points $\Gamma$, 0.58-0.58] and [X, $0.01$] in the 155v electronic structure (Fig. 4, panel b). The wavefunction is represented as its contribution to the electronic charge density at an isovalue of 0.003 a.u.
Unless otherwise specified, all Seebeck coefficients and electronic conductances discussed below correspond to $T=400~\text{K}$. As seen in previous theoretical literature, the measurable electronic parameters of thermoelectricity—namely the Seebeck coefficient and electric conductivity—can be calculated from DFT, yielding satisfactory agreement with experimental values even without introducing any corrections beyond a rigid shift of the bands [20, 54, 24].
Within our theoretical approach, we obtain quantitative results by consistently comparing coherent systems, e.g. a system containing an impurity with the corresponding pristine system of the same dimensionality. When we compare energy dependent functions—such as transmission, Seebeck coefficient and conductance of corresponding systems (namely in Figures 6, 8, 9, 10, 11, 12, 13)—we shift them on the energy scale to ensure structural correspondence. Specifically, this is best achieved in (1×)5×5+s systems when the transmittances of the quantum channels associated with the highest-energy valence bands are aligned, and in (3×)3×3+s systems when the central values of the fundamental gaps are matched at the same energy.
@@
Figure 6: Panel a: Calculated electronic conductances at 0 K as functions of the chemical potential of (1x)5x5+s ScN NW systems: pure (155); BT defected system with a vacancy replacing the central N atom (155v): relaxed and unrelaxed; and a BT defected system in which a glide-type defect is contiguous to the N site vacancy and perpendicular to NW (155v+gl). Panel b: Calculated Seebeck coefficients as functions of the chemical potential at 400 K for the three systems 155, BT 155v (relaxed) and BT 155v+gl. Vertical dashed lines represent the calculated Fermi levels of the systems.
@@
In Figures 9 and 10 we present the resulting transmission functions and Seebeck coefficients profiles for the point-defected BT systems with central O impurities in place of N sites. The results of our calculations show that this kind of impurities have negligible effects on the profiles of Seebeck coefficients when compared to the effects of vacancies on N sites. This holds true whether the O impurities are isolated ( [^3] O) or packed in contiguous positions (155O).
In Figures 6 and 9, alongside the pristine and BT single‐point‐defected (1x)5x5+s systems, we also present the calculated transmission functions and Seebeck coefficient functions of systems combining—in contiguous positions and BT configuration—glide‐type and point defects (155v+gl and 155O+gl). We observe that the presence of a glide‐type defect produces a pronounced decrease in the Seebeck coefficient profile in both cases, comparable to that induced by multiple vacancies at contiguous N sites. This allows us to understand that the effects related to vacancies association (v+v) are primarily electronic in nature when compared to the effects of oxygen associations (O+O or O+gl), since they manifest similarly whether or not the vacancy is paired with a glide‐type defect, whereas the impact of an oxygen impurity adjacent to a glide‐type defect (O+gl) is predominantly structural in nature when compared to v+v, as its profile reduction of the Seebeck function disappears in the absence of the glide‐type defect.
The v+v configurations lead to a substantial increase in electronic conductance. We computed different electric conductivity functions $\sigma(\mu)$ of ScN from the different conductance functions of our NW models, introducing a rigid shift of the conduction bands to reproduce the experimental 0.9 eV gap:

### v3

# Thermoelectric power factors of defective scandium nitride nanostructures from first principles

@@ Abstract
The thermoelectric properties of scandium nitride are strongly influenced by structural and electronic factors arising from defects and impurities. Nevertheless, the mechanisms by which these microscopic features affect transport are not yet fully understood. Experiments show a large variability in the electronic transport properties, with a strong dependence on the experimental conditions, and attempts to improve thermoelectric efficiency often lead to conflicting effects. In this work, we employ the Landauer approach to analyze the effects of different kinds of structural defects and impurities on electronic transport in scandium nitride. This approach allows us to relate the transport mechanisms to the structural and electronic modifications introduced in the lattice, with atomistic resolution. In light of these new insights, we propose a rationale relating part of the experimental variability to its microscopic origin.
@@ I Introduction
Appropriate preparation techniques for ScN thin films can also significantly reduce their thermal conductivities, as observed with lithium (Li) ions irradiation during plasma-assisted molecular beam epitaxy [59], or by implanting argon (Ar) or atoms or Helium (He) ions into ScN thin films during or after reactive magnetron sputtering deposition using an ion beam implantation technique [11, 10]. These procedures lead to a strong decrease in the in-plane thermal conductivity of the resulting films, reaching respectively 7 and 3 $\mathrm{W/(m\cdot K)}$ at RT. Recently, a similar reduction has been reported on ScN thin films prepared by reactive magnetron sputtering, with periodic interruptions of the cathodic potential during the deposition procedure [48]. However, in addition to their impact on thermal transport, these techniques also affect electronic transport, leading to opposing effects—regarding thermoelectric efficiency—between the Seebeck coefficient and electrical conductivity. Specifically, they induce a significant increase—in absolute terms—of the Seebeck coefficient while causing a strong reduction in electrical conductivity. The total effect on the $zT$ value is slightly positive, with an increase in the thermoelectric figure of merit primarily driven by the substantial reduction in thermal conductivity, despite a decrease in the thermoelectric power factor [59, 11, 48, 10].
Theoretical and experimental studies have correlated the transport properties of ScN with its structural defects. The presence of chemical impurities and lattice imperfections has a strong impact on the physical properties of ScN even under room-temperature sputtering deposition, and an even larger one under standard high-temperature deposition conditions [16]. Impurities and lattice vacancies have consequently been identified as key factors for thermoelectricity [33, 43, 52, 61, 37, 57, 17]. However, the mechanisms through which these microscopic features regulate the thermoelectric properties remain not fully investigated. Understanding these microscopic regulation mechanisms is crucial for developing further optimization strategies capable of selectively tuning individual parameters without negatively affecting others.
In this work, we aim to achieve a deeper understanding of the microscopic mechanisms regulating electronic transport—and, consequently, thermoelectricity—in ScN. We aim to do so by modeling the electronic transport in ScN nanowires (NW) and associating it with structural lattice defects, whose formation in the ScN lattice can be hypothesized under the described experimental conditions [12, 59, 11, 48, 10]. We find general conclusions applicable to ScN lattices regardless of their dimensionality. Specifically, we study point defects such as the inclusion of oxygen (O) impurities and atomic vacancies in place of the lattice nitrogen (N) atoms. Furthermore, we examine the effects of association of multiple defects in contiguous lattice positions and the effects of associacion with stacking faults.
@@ II.1 The Landauer approach to quantum transport
$\Sigma_{\{s,d\}}(E)$ and $\Sigma_{\{s,d\}}^{\dagger}(E)$ are the retarded and advanced self-energies of the leads (source or drain), which contain the retarded and advanced Green’s functions of the leads and the coupling matrices between the central conductor and the leads. Their rigorous expression can be found e.g., in Refs. 51, 19, 14. The retarded and advanced Green’s functions of the central conductor, $\mathcal{G}_{c}^{r}(E)$ and $\mathcal{G}_{c}^{a}(E)$, describe the propagation of the electrons forwards and backwards in time in this section of the system. These can be straightforwardly calculated once knowing the real-space Hamiltonian of the system and the Green’s functions of the semi-infinite leads [51, 14, 19].
In this work, we calculate the Bloch space Hamiltonians of the systems using plane waves Density Functional Theory [35] and the real-space Hamiltonians with a subsequent projection and filtering procedure [3]. The Green’s functions of the semi-infinite leads are then calculated dividing the real-space representation of the system into a series of principal layers orthogonal to the transport direction (see Fig. 1) and then using the iterative Lopez-Sancho et al. procedure [64, 15].
@@ II.2 Structural models
As mentioned, ScN has a face-centered cubic (fcc) rocksalt structure with a calculated lattice constant of 4.530 Å. The primitive cell contains 2 atoms per cell, while it is possible to define a conventional simple cubic (sc) cell using a larger crystal unit, containing 8 atoms per cell. The axes of the conventional sc cell are aligned along the Sc–N bonds directions.
We chose to study the defective structures by embedding them in NW models. These nanostructures, being finite along the directions transverse to transport, allow us to obtain quantized, energy-dependent transmittance functions that can be directly compared with the electronic band structures of the systems, revealing important insights into the electronic transport mechanism that would not be accessible with 3D models. Using periodic boundary conditions (PBC), we built the NW models along one of the Sc–N direction. The NWs were constructed by inserting sufficient space in the directions orthogonal to their length and replicating half of the surface layers of the obtained NW on the opposite sides (these additional layers are represented by +s in the symbolism that we introduce to denote the NWs’geometries). This operation was done to make the cross-section of the NWs centrosymmetric. Two different periodicities were employed, corresponding to either one or three conventional cells along the NW direction, and different cross-section areas: (1x)1x1+s, (1x)3x3+s, (1x)5x5+s, (3x)3x3+s. These symbols describe the models’ structures: in parentheses the length of the periodic unit—expressed in conventional cells—then the cross-sectional area, and the symbol +s for additional layers.
@@
Figure 2: Schematic representations of (a) a stacking fault, (b) a point defect, and (c) a combined configuration of a planar stacking fault and a point defect.
The possible structural defects that may arise in the ScN lattice can be classified according to their symmetry into point defects and planar faults. Combinations of these two types of defects can also occur in spatially adjacent positions (Figure 2).
We consider stacking faults (sf) perpendicular to the transport direction, in order to maximize the effects of the structural mismatches and observe them more clearly. We modeled these planar defects in the (1x)5x5+s structures, associating them with point defects. After obtaining the relaxed NW structures with point defects, we replaced a relaxed atomic layer adjacent to the defect and perpendicular to the NW length, with a layer from the unrelaxed structure, centered along the same NW axis.
Several types of point-type defects can be accounted: vacancies and substitutions of either Sc or N atoms, or clusters of multiple such defects. These defects can be ranked based on their relevance by calculating their formation enthalpy, revealing that substitutional N defects are the most energetically favorable and thus the most common in experimental samples [7].
Concerning the chemical nature of defects, under magnetron sputtering deposition or molecular-beam epitaxy experimental conditions [34, 12, 59, 11, 48], the presence of oxygen impurities is likely unavoidable, even in ultra-high-vacuum (UHV) systems [46, 47, 18]. We studied substitutional defects where N sites are replaced either by structural vacancies or by oxygen atoms, and association of these structures with contiguous point defects or planar mismatches. We coherently compared each case to models of pure ScN nanowires of the same size.
@@ II.3 Detailed description
The atomic positions in systems with structural point defects were obtained through structural relaxation calculations performed using the Broyden–Fletcher–Goldfarb–Shanno (BFGS) algorithm [30]. These structural relaxations were carried out on bulk systems, constructed with PBC. We introduced the structural defect at the center of supercells of size 3x3x3 and 1x5x5 conventional cells. For the structural relaxation calculations, the eigenvalues were calculated on uniform 3x3x3 k-points grids spanning the whole first Brillouin zone of the systems, exploiting space-group symmetries where possible; we employed a PW basis set defined by a cutoff energy of $40\,\text{Ry}$; the electronic charge density was recalculated after each diagonalization using a larger set of plane waves, defined by a cutoff energy of $320\,\text{Ry}$, before proceeding to the next step in the self-consistent field (SCF) cycle. The convergence thresholds were set to $10^{-6}$ a.u. for the total energy and $10^{-5}$ a.u. for the forces.
@@ III Results and discussion
III.1 Electronic band structures
In Figure 3, we present the calculated electronic band structures of pure ScN represented along high-symmetry lines of the primitive (p) face-centered cubic (fcc) lattice—containing two atoms per unit cell—and the conventional (c) simple cubic (sc) lattice—containing eight atoms per cell  [^1] The electronic band structures of the bulk three-dimensional systems were calculated as described in section II.3, using an 81×81×81 k-point grid spanning the first Brillouin zone of the primitive cell and a 21×21×21 k-point grid for the simple cubic conventional cell, exploiting space-group symmetries where possible to reduce the computational effort.. The indirect band gap $\Gamma\text{--}X$]p observed in the fcc representation appears as a direct gap at the $\Gamma$ point in the sc representation, due to band folding at the $\Gamma$ point. Theoretical literature reports an indirect $\Gamma\text{--}X$]p band gap of $\sim 0.9$ eV for pure ScN, calculated using quasiparticle corrections [58, 39], exact-exchange approach [24] or hybrid functionals [20, 50]. These theoretical values are in good agreement with experimental literature [21, 56, 24]. This indirect band gap, however, is not captured by DFT, due to the well-known underestimation of the fundamental gap. In either the local density approximation (LDA) or GGA (Figure 3), it is predicted to be zero [6]. Apart from this known issue, the electronic structure computed using DFT does not exhibit significant topological or analytic variations in the shape of the bands near to the Fermi energy when compared to those obtained using quasiparticle corrections, exact exchange or hybrid functionals [58, 33].
@@ III.1 Electronic band structures
In Figure 4, we compare the electronic band structures of the defected (1×)5×5+s systems with central oxygen atoms and vacancies. While the presence of oxygen only leads to an $n$-type doping effect, with a rigid shift of the Fermi level without significantly distorting the band topology, the system with vacancies, in addition to a similar effect, exhibits also a strong distortion in the shape of a single band—whose real-space representation is shown in Figure 5—to the point that it crosses the Fermi level. This, in turn, affects the profile of the electronic transmission function (Figure 6), with subsequent effect on the Seebeck coefficient. In order to exclude the possibility that this feature results from nanostructuring effects in the NW model, we calculated the electronic band structure of the same systems by removing the additional +s layers and reducing the cell dimensions along the directions transverse to the NW length, thereby restoring the 3D periodicity. As a result, we confirm that this feature is also present in this case, demonstrating that it is not a consequence of the reduced dimensionality.
@@
Figure 5: A section of the 155v NW system. In gray: real-space representation at the $\Gamma$-point of the wavefunction associated with the electronic state whose band connects the points ($\Gamma$, 0.58-0.58) and (X, $0.01$) in the 155v electronic structure (Fig. 4, panel b). The wavefunction is represented as its contribution to the electronic charge density at an isovalue of 0.003 a.u., using the XCrySDen visualization tool [36].
III.2 Seebeck coefficients
As seen in previous theoretical literature, the measurable electronic parameters of thermoelectricity—namely the Seebeck coefficient and electric conductivity—can be calculated from DFT, yielding satisfactory agreement with experimental values even without introducing any corrections beyond a rigid shift of the bands [12, 6, 48].
Within our theoretical approach, we obtain quantitative results by consistently comparing coherent systems, e.g. a system containing an impurity with the corresponding pristine system of the same dimensionality. When we compare energy dependent functions—such as transmission, Seebeck coefficient and conductance of corresponding systems (namely in Figures 6, 8-13)—we shift them on the energy scale to ensure structural correspondence. Specifically, this is best achieved in (1×)5×5+s systems when the transmittances of the quantum channels associated with the highest-energy valence bands are aligned, and in (3×)3×3+s systems when the central values of the fundamental gaps are matched at the same energy.
@@ III.2 Seebeck coefficients
Figure 6: Panel a: Calculated electronic conductances at 0 K as functions of the chemical potential of (1x)5x5+s ScN NW systems: pure (155); BT defected system with a vacancy replacing the central N atom (155v): relaxed and unrelaxed; and a BT defected system in which a stacking fault is contiguous to the N site vacancy and perpendicular to NW (155v+sf). Panel b: Calculated Seebeck coefficients as functions of the chemical potential at 400 K for the three systems 155, BT 155v (relaxed) and BT 155v+sf. Vertical dashed lines represent the calculated Fermi levels of the systems.
@@
Figure 9: Comparison of the calculated conductances at 0 K as functions of the chemical potential for the 3×(3×3)+s ScN NW models: pristine BT 333 and TL model composed by 333vv as central conductor and pristine 333 as leads (333-333vv-333).
In Figures 10 and 11 we present the resulting transmission functions and Seebeck coefficients profiles for the point-defected BT systems with central O impurities in place of N sites. The results of our calculations show that this kind of impurities have negligible effects on the profiles of Seebeck coefficients when compared to the effects of vacancies on N sites. This holds true whether the O impurities are isolated ( [^3] O) or packed in contiguous positions (155O).
In Figures 6 and 10, alongside the pristine and BT single‐point‐defected (1x)5x5+s systems, we also present the calculated transmission functions and Seebeck coefficient functions of systems combining—in contiguous positions and BT configuration—stacking faults and point defects (155v+sf and 155O+sf). We observe that the presence of a stacking mismatch produces a pronounced decrease in the Seebeck coefficient profile in both cases, comparable to that induced by multiple vacancies at contiguous N sites. This allows us to understand that the effects related to vacancies association (v+v) are primarily electronic in nature when compared to the effects of oxygen associations (O+O or O+sf), since they manifest similarly whether or not the vacancy is paired with a stacking fault, whereas the impact of an oxygen impurity adjacent to a planar defect (O+sf) is predominantly structural in nature when compared to v+v, as its profile reduction of the Seebeck function disappears in the absence of the stacking mismatch.
Figure 10: Panel a: Calculated electronic conductances at 0 K as functions of the chemical potential of (1x)5x5+s ScN NW systems: pure (155); BT defected system with an O atom replacing the central N site (155O): relaxed and unrelaxed; and a BT defected system in which a stacking fault is contiguous to the central O impurity and perpendicular to NW (155O+sf). Panel b: Calculated Seebeck coefficients as functions of the chemical potential at 400 K for the three systems 155, BT 155O (relaxed) and BT 155O+sf. Vertical dashed lines represent the calculated Fermi levels of the systems.
In Figure 12, we report the calculated Seebeck functions for selected BT 3×(3×3)+s ScN NW systems. These functions were obtained after a rigid shift of the conduction bands in order to reproduce the 0.9 eV experimental gap. It can be observed that this procedure yields numerically reasonable results when compared with experimental values. However, this approximation overestimates the distance between the calculated chemical potential and the peak values of the Seebeck functions. Indeed, reducing the cross-section of the NW models leads to increasingly higher absolute peak values in the calculated Seebeck functions, and similar effects are observed when a rigid shift is applied to the energy bands.
To summarize, we have identified two notable categories of defect structures that produce a significant effect on the shape of the Seebeck functions: v+v and O+sf. These two categories of defects might be responsible for part of the experimental variability in the Seebeck coefficient reported in the experimental literature [12, 59, 11, 48, 10] and will be further investigated in the following subsection.
III.3 Electronic conductivities
The band structure of the 155v model (Fig. 4) suggests that the presence of v+v configurations can lead to a substantial increase in electronic conductance. After introducing — using the same procedure described above — a rigid shift of the conduction bands to reproduce the experimental 0.9 eV band gap, we computed electric conductivities $\sigma(\mu)$ as functions of the chemical potentials from the different conductance functions G of our NW models: